OpenStudy (anonymous):

Simplify. Please show all work. 12/1/4+2/5-2^3

jhonyy9 (jhonyy9):

12/1/4 =12 * 4/1 = 12*4 =48 2/5 -2^3 so if -2^3 is denominator too this is equal 2/5-8 =2/-3 so and the result will be 48-2/3 = (144-2)/3 = 142/3